OpenStudy (anonymous):

I have a recipe that they are adding a 1/2 cup of flour to. I am suppose to redo the recipe but I dont know if I'm doing it right. Start off with 1/2 brown sugar, 1 1/2 c. oats, 1 1/2 c. flour, 1/2 Tbsp baking power, 1/2 bsp. baking soda, 1/2 tsp salt, 2 eggs, 1 c. applesauce, 4 Tbsp. vegetable oil. Now I have 2 add the 1/2 flour and the rest of the recipe has to increase and I dont know what the answer would be?

OpenStudy (anonymous):

>ou got 1 1/2 = 1.5 c flour. Addin 1/2 your're increasing it by one third or 33.33% to 2 c of flour. Multiply each compound of the recipe with 1.33 and you get the right amount for eatch recipe

OpenStudy (anonymous):

*You got

OpenStudy (anonymous):

\[\log_{3}\left( 3^{x+2} \right) = \log_{3}4^{x} +1\] \[x+2=xlog_{3}4 + 1\]

OpenStudy (anonymous):

What do u want to prove?? The task is just simple percentage calculation.